OpenStudy (anonymous):

50^2+60^2=N√625 please show steps

OpenStudy (anonymous):

N√(625) = 50^(2) + 60^(2) N√(625) = 6100 N = 6100 / √(625) // Divide by √(625) on both sides N = 6100 / 25 // √(625) = 25 N = 244

OpenStudy (anonymous):

Words following "//" are comments.